### v2

@@ A.3 Extension to Structured Non-Boolean Matrices
Definition A.14 (Pollard Pseudodimension).
For a family of functions $F$ mapping $V$ to $\{0,\dots,\ell-1\}$, the Pollard pseudodimension of $\mathcal{F}$, denoted by $\mathrm{Pdim}(\mathcal{F})$, is the largest integer $d$ such that there exists $x_{1},\dots,x_{d}\in V$ and thresholds $y_{1},\dots,y_{d}\in[\ell]$ such that for any $d$-length bit vector $\vec{b}=(b_{1},\dots,b_{d})\in\{0,1\}^{d}$, there exists an $f\in\mathcal{F}$ such that $\forall i,f(x_{i})\leq y_{i}\iff b_{i}=1$.
As an example, consider an integer matrix and split each row into the different thresholds.
$\begin{bmatrix}1&2&4&2\end{bmatrix}\to\begin{bmatrix}0&0&0&0\\ 1&0&0&0\\ 1&1&0&1\\ 1&1&1&1\end{bmatrix}$
Then, the right matrix has VC dimension $d$, if and only if the left matrix has Pollard pseudodimension $d$.
Definition A.15 (Multiball vectors).
Let $v\in V$ be a vertex of the graph $G$. Let $r$ be a real number and $\Delta$ be a set of distances $-\infty=\delta_{0}<\delta_{1}<\dots<\delta_{\ell-1}<\delta_{\ell}=\infty$. Define the multiball vectors as follows:
$\vec{\mathrm{MB}}(v,r,\Delta)=(y_{u})_{u\in V}\text{ where }y_{u}\in[\ell]\text{ is the smallest integer such that }u\in\vec{B}(v,r+\delta_{yu}).$
@@
Theorem A.16 (Karczmarz and Zheng (2024)).
@@
For a $K_{h}$-minor-free graph $G$, for any set $\Delta$ of $\ell$ distances, the set of multiball vectors $\vec{\mathrm{MB}}_{G,\Delta}$ has pseudodimension at most $h-1$.
Lemma A.17 (Generalized Sauer-Shelah lemma).
Let $\mathcal{F}$ be a family of functions mapping a ground set with $n$ elements to $[\ell]$. If the pseudodimension of $\mathcal{F}$ is at most $d$, then $|\mathcal{F}|\leq O(n^{d}\ell^{d})$.
Lemma A.18.
Let $\mathcal{F}$ be a family of functions mapping a ground set with $n$ elements to $[\ell]$. Let $\mathcal{F}_{s}=\mathcal{F}\cap[\ell]^{S}$ be a restriction of $\mathcal{F}$ to $S$. Then, $\mathrm{Pdim}(\mathcal{F}_{S})\leq\mathrm{Pdim}(\mathcal{F})$.
Corollary A.19.
Let $\mathcal{F}$ be a family of functions mapping a ground set with $n$ elements to $[\ell]$, and let $S$ be any subset of the ground set of $\mathcal{F}$. Suppose $\mathrm{Pdim}(\mathcal{F})=d$. Then, $|\mathcal{F}_{S}|=O(|S|^{d}\ell^{d})$.
Definition A.20.
For a real matrix $\mathbf{M}$, define the pseudo-dimension of $\mathbf{M}$, denoted by $\mathrm{Pdim}(\mathbf{M})$ by thinking of the rows of $\mathbf{M}$ as functions and taking the pseudo-dimension of the resulting class of functions. Specifically, if $\mathbf{M}$ is an $m\times n$ matrix, for each $i\in\{1,\dots,m\}$, define $f_{\mathbf{M},i}:\{1,\dots,n\}\to\mathbb{R}$ by $f_{\mathbf{M},i}(j)=\mathbf{M}_{i,j}$ and let $\mathrm{Pdim}(\mathbf{M})=\mathrm{Pdim}(\{f_{\mathbf{M},i}:i\in\{1,\dots,m\}\})$.
Lemma A.21.
Given any matrix $\mathbf{M}\in\mathbb{Z}^{n\times n}$, construct the Boolean matrix $\mathbf{N}\in\{0,1\}^{n\times n}$ such that $\mathbf{N}_{i,j}=\mathbbm{1}\{\mathbf{M}_{i,j}\geq 0\}$ for $i,j\in[n]$. Then, $\mathrm{VC}(\mathbf{N})\leq\mathrm{Pdim}(\mathbf{M})$.
Proof A.22.
Given a matrix $\mathbf{M}$, we have threshold functions $f_{\mathbf{M},i}$ such that $f_{\mathbf{M},i}(j)=\mathbf{M}_{i,j}$ and can use this this to define a set family
=$\mathcal{F}=\{f_{\mathbf{M},i}:\{1,\dots,n\}\to\mathbb{R}|i=1,\dots,m\},$
and $\mathrm{Pdim}(\mathbf{M})=\mathrm{Pdim}(\mathcal{F})$.
By definition, $\mathrm{Pdim}(\mathbf{M})$ is the largest integer $d$ such that $\exists x_{1},\dots,x_{d}\in\{0,\dots,\ell-1\}$ and thresholds $y_{1},\dots,y_{d}\in[\ell]$ such that for any vector $b_{1},\dots,b_{d}\in\{-1,1\}^{d}$, there is an $f\in\mathcal{F}$ such that $\forall i,\mathrm{sign}(f(x_{i})-y_{i})=b_{i}$.
Define the following $0/1$ profile of matrix $\mathbf{M}$. For each row $i$ of $\mathbf{M}$, define the binary function
$g_{\mathbf{M}.i}:\{1,\dots,n\}\to\{0,1\}$
by
$g_{\mathbf{M},i}(j)=\mathbbm{1}\{f_{\mathbf{M},i}(j)\geq 0\}\coloneqq\mathbbm{1}\{\mathbf{M}_{i,j}\geq 0\},$
and use this to define the set family
$\mathcal{G}=\{g_{\mathbf{M},i}:i=1,\dots,m\}.$
We want to show that if $\mathrm{Pdim}(\mathbf{M})$ is low ($\mathrm{Pdim}(\mathbf{M})=d$), then the VC-dimension of $\mathcal{G}$ is at most $d$. So, by way of contradiction, suppose $\mathrm{Pdim}(\mathcal{F})=d$ and $\mathrm{VC}(\mathcal{G})\geq d+1$.
Then, there must exist $S\subset\{1,\dots,n\}$ with $|S|=d+1$ that is shattered by $\mathcal{G}$. So, for every $0/1$ labeling of $c:S\to\{0,1\}$, there is some row $i$ for which $g_{\mathbf{M},i}(j)=c(j),\forall j\in\mathcal{S}$.
Also, $\exists x_{1},\dots,x_{d}\in\{0,\dots,\ell-1\}$ and thresholds $y_{1},\dots,y_{d}\in[\ell]$ such that for any vector $b_{1},\dots,b_{d}\in\{-1,1\}^{d}$, there is an $f\in\mathcal{F}$ such that $\forall i\in[d],\mathrm{sign}(f(x_{i})-y_{i})=b_{i}$.
Also, note that the condition
$g_{\mathbf{M},i}(j)=\mathbbm{1}\{f_{\mathbf{M},i}(j)\geq 0\}=c(j)$
is equivalent to
$\mathrm{sign}(f_{\mathbf{M},i}(j)-0)=b^{*}_{j},\text{ for }b^{*}_{j}=\begin{cases}+1,&c(j)=1,\\ -1,&c(j)=0\end{cases}$
but this just means that $S$ is pseudo-shattered by the class $\mathcal{F}$, where the thresholds $y_{1},\dots,y_{d}=0$ for all $j\in S$. So, this implies $\mathrm{Pdim}(\mathbf{M})=d+1$, which is a contradiction. So, $\mathrm{VC}(\mathcal{G})\leq\mathrm{Pdim}(\mathcal{F})$.
Lemma A.23.
Suppose that the Pollard pseudodimension of the matrix $\mathbf{M}\in\mathbb{R}^{m\times n}$ is $d$, and that the number of thresholds is $A$. Then, after an $\widetilde{O}(Amn)$ time preprocessing, there is a data structure that, upon receiving a vector $v\in\mathbb{R}^{n}$, can return $\mathbf{M}v$ in time $O(Amn^{1-1/d}+An)$.
Proof A.24.
Given the matrix $\mathbf{M}\in\mathbb{R}^{n\times n}$ with $A$ thresholds $\tau_{1},\tau_{2},\dots,\tau_{A}\in\mathbb{R}^{A}$, construct the Boolean matrix $\mathbf{N}\in\{0,1\}^{Am\times n}$ as follows: for row $i\in[n]$ of $\mathbf{M}$, define the block $\mathcal{B}^{(i)}\coloneqq\mathcal{B}(\mathbf{M}_{i})\in\{0,1\}^{A\times n}$ by
$\mathcal{B}(\mathbf{M}_{i})=\begin{bmatrix}\mathbbm{1}_{\{\mathbf{M}_{i,1}\geq\tau_{1}\}}&\mathbbm{1}_{\{\mathbf{M}_{i,2}\geq\tau_{1}\}}&\dots&\mathbbm{1}_{\{\mathbf{M}_{i,n-1}\geq\tau_{1}\}}&\mathbbm{1}_{\{\mathbf{M}_{i,n}\geq\tau_{1}\}}\\ \mathbbm{1}_{\{\mathbf{M}_{i,1}\geq\tau_{2}\}}&\mathbbm{1}_{\{\mathbf{M}_{i,2}\geq\tau_{2}\}}&\dots&\mathbbm{1}_{\{\mathbf{M}_{i,n-1}\geq\tau_{2}\}}&\mathbbm{1}_{\{\mathbf{M}_{i,n}\geq\tau_{2}\}}\\ &&\ddots&&\\ \mathbbm{1}_{\{\mathbf{M}_{i,1}\geq\tau_{A}\}}&\mathbbm{1}_{\{\mathbf{M}_{i,2}\geq\tau_{A}\}}&\dots&\mathbbm{1}_{\{\mathbf{M}_{i,n-1}\geq\tau_{A}\}}&\mathbbm{1}_{\{\mathbf{M}_{i,n}\geq\tau_{A}\}}\end{bmatrix}$
Then, matrix $\mathbf{N}\in\{0,1\}^{Am\times n}$ is given by
$\mathbf{N}=\begin{bmatrix}\mathcal{B}^{(1)}\\ \mathcal{B}^{(2)}\\ \vdots\\ \mathcal{B}^{(m)}\end{bmatrix}$
By lemma A.21, the VC-dimension of $\mathbf{N}$ is at most the Pollard pseudodimension of $\mathbf{M}$, given by $d$. Constructing $\mathbf{N}$ takes $O(mnA)$ time. Then, using theorem 1.1, we can preprocess $\mathbf{N}$ in $\widetilde{O}(Amn)$ time to support queries that return $\mathbf{N}v$ for any vector $v\in\mathbb{R}^{n}$ in time $\widetilde{O}(Amn^{1-1/d}+n)$. To convert $\mathbf{N}v\in\mathbb{R}^{Am}$ to $\mathbf{M}v\in\mathbb{R}^{m}$, iterate over each of the $m$ sets of $A$ entries of $\mathbf{N}v$. For $i\in[m]$, denote the $i$’th set of $A$ entries by $x^{(1)}_{i},x^{(2)}_{i},\dots,x_{i}^{(A)}$, where $x_{i}^{(j)}\coloneqq(\mathbf{M}v)_{A(i-1)+j}$. Then, defining $y_{i}$ as the linear combination of the threshold maps with the outputs $x_{i}^{(j)}$, we have $y_{i}\coloneqq\sum_{j=1}^{A}x_{i}^{(A)}\tau_{i}$. Therefore, $\mathbf{N}v$ is given by
$\mathbf{N}v=\begin{bmatrix}y_{1}\\ y_{2}\\ \vdots\\ y_{m}\end{bmatrix},$
which proves the claim.
Appendix B Applications
@@
In this section, we consider some applications of our OMv result in Theorem 1.2. These applications have quadratic $\Omega(n^{2})$-time lower bounds, conditional on the OMv conjecture. We show that on structured graph this lower bound can be beaten.
B.1 High-accuracy Dynamic Laplacian Solver
See 1.4
@@ B.1 High-accuracy Dynamic Laplacian Solver
Definition B.1 ($(1+\epsilon)$-spectral sparsifiers).
@@
Then, a result from Abraham et al. (2016) provides a construction for a dynamic spectral sparsifier under edge deletions and insertions with polylogarithmic amortized update time:
Theorem B.2 (Theorem 4.1 of Abraham et al. (2016)).
@@
Lemma B.3 (Kyng and Sachdeva (2016)).
@@
Via Theorem 1.2 we have $U(n)=\widetilde{O}(n)$ and $Q(n)=\widetilde{O}(n^{2-1/d})$, thus obtaining Theorem 1.4.
Proof B.5.
We run two separate data structures. The first is a dynamic spectral sparsifier Theorem B.2 which maintains a spectral sparsifier $H\approx\mathbf{L}_{G}$ with error $\epsilon=1/2$. We also run the dynamic OMv data structure for matrix $\mathbf{A}$ (adjacency matrix).
@@
When given a vector $b\in\mathbb{R}^{V}$ and error parameter $\epsilon>0$, we use $\mathbf{H}^{-1}$ as preconditioner for the linear system $\mathbf{L}_{G}x=b$. We can multiply with $\mathbf{H}^{-1}$ by running a static Laplacian system solver Lemma B.3 which takes $\widetilde{O}(n)$ time per product as $\mathbf{H}$ has $\widetilde{O}(n)$ edges. In particular, we let
@@
B.2 Effective Resistance
The following Theorem 1.5 is a corollary of the dynamic Laplacian solver from Theorem 1.4.
See 1.5
Proof B.6.
@@ B.2 Effective Resistance
This can be computed via the dynamic Laplacian system solver in Theorem 1.4.
B.3 Triangle Detection
See 1.6
The result follows form the following lemma for $U(n)=\widetilde{O}(n)$ and $Q(n)=\widetilde{O}(n^{2-1/d})$ from Theorem 1.2.
Lemma B.7.
@@ B.3 Triangle Detection
B.4 Single-Source Shortest Paths and k-Center
Lemma B.9 (van den Brand et al. (2022)).
@@ B.4 Single-Source Shortest Paths and k-Center
A query operation which for any source node, given during the query, returns the $d$-bounded single source distances in $O(Q(n,d))$. That is, return for each $v\in V$ the distance if it is at most $d$, or $\infty$ if the distance is $>d$.
@@
See 1.7
Proof B.10.
We first describe how to obtain a dynamic distance oracle that yields distances up to $1/\epsilon$. This is done by running the dynamic OMv data structure of Theorem 1.2 on the adjacency matrix of $G$. Then single source distances up to $1/\epsilon$ from source vertex $u$ can be computed by repeatedly multiplying $w\leftarrow\mathbf{A}w$ for initial $w=e_{u}$, i.e., we compute $\mathbf{A}^{i}e_{u}$ for $i=1,2,...,1/\epsilon$. The smallest $i$ with $(\mathbf{A}^{i}e_{u})_{v}\neq 0$ is the smallest number of steps to reach from $u$ to $v$, i.e., the distance between the vertices.
This data structure internally constructs the a $\Delta$-labeled spanning tree of the boolean matrix $\mathbf{A}$, where the weight of the tree gives us the time complexity per query.
Hence, while we do not know the corrupted VC-dimension $d$, we do know how much time each query is going to take.
So when running Lemma B.9 on the above distance data structure, we can pick $k=n/\sqrt{T}$ where $T\geq n$ is the weight of the spanning tree. Thus, we get the update time
@@
which proves the claim.
See 1.8
This result is a corollary of Theorem 1.7, as $k$-center can be reduced to computing $k$-single source distances.
Lemma B.11 (Theorem 7.3 by Cruciani et al. (2024)).
@@
Appendix C Characterization of matrices with constant VC-dimension
In this section, we provide a characterization of matrices with constant VC-dimension. Particularly, we prove section 1 and state a number of examples of other constant VC-dimension matrices.

### v3

@@ B.4 Extension to Structured Non-Boolean Matrices
Definition B.11 (Pollard Pseudodimension).
For a family of functions $\mathcal{F}$ mapping $X$ to $Y$, the Pollard pseudodimension of $\mathcal{F}$, denoted by $\mathrm{Pdim}(\mathcal{F})$, is the VC dimension of the set system $(X\times Y,\{\{(x,y)\mid f(x)\geq y\}\mid f\in\mathcal{F}\})$.
In the setting of matrices $\mathbf{M}\in\mathbb{R}^{m\times n}$, we interpret each row as a function $f_{\mathbf{M},i}:x\mapsto\mathbf{M}_{i,x}$. In particular, when $\mathbf{M}$ has Pollard pseudodimension $d$, then that means the set system $([n]\times\mathbb{R},\{(j,y)\mid\mathbf{M}_{i,j}\geq y\}\mid i\in[m]\})$ has VC-dimension $d$. Restricting the groundset $n]\times\mathbb{R}$ to $\{(j,y)\mid\exists i:\mathbf{M}_{i,j}=y\}$ does not increase the VC-dimension. This latter set system can be interpreted as the following Boolean matrix $\mathbf{N}$:
Given a matrix $\mathbf{M}\in\mathbb{R}^{m\times n}$ where each column has at most $T$ distinct values, consider the following construction of an $\mathbf{N}\in\{0,1\}^{m\times nT}$ matrix. Let $\mathbf{Y}\in\mathbb{R}^{T\times n}$ be the distinct values per column of $\mathbf{M}$ (we may duplicate values if a column has <T<T distinct values). Assume the entries in each column of $\mathbf{Y}$ are sorted in increasing order. Let $\mathbf{N}_{i,jT+k}=\mathbbm{1}_{\mathbf{M}_{i,j}\geq\mathbf{Y}_{i,k}}$.
For example, the following $\mathbf{M}$ has at most $T=4$ distinct values per column:
$\mathbf{M}\coloneqq\left[\begin{array}[]{c|c}1&-1\\ 4&-2\\ 2&4\\ 5&3\\ 4&3\end{array}\right]\to\left[\begin{array}[]{cccc|cccc}1&0&0&0&1&0&0&0\\ 1&1&0&0&1&1&0&0\\ 1&1&1&1&1&1&1&1\\ 1&1&1&0&1&1&1&0\\ 1&1&0&0&1&1&1&0\end{array}\right]\eqqcolon\mathbf{N},\quad\mathbf{Y}:=\begin{bmatrix}1&-2\\ 2&-1\\ 4&3\\ 5&4\end{bmatrix}$
Since matrix $\mathbf{N}$ is Boolean and of VC-dimension at most $d$, we can run our fast matrix-vector product data structures on it (e.g., Theorem˜1.3).
Proof of Theorem˜B.9.
Given $\mathbf{M}\in\mathbb{R}^{m\times n}$ we construct the Boolean matrix $\mathbf{N}\in\{0,1\}^{m\times nT}$ as previously described.
We initialize Theorem˜1.3 on $\mathbf{N}$ which takes $\widetilde{O}(Tmn)$ time. Multiplying a vector $v^{\prime}\in\mathbb{R}^{nT}$ with $\mathbf{N}$ takes $\widetilde{O}(Tnm^{1-1/d}+m)$ time. We are left with explaining how to compute $\mathbf{M}v$ via $\mathbf{N}v^{\prime}$.
We have
$\mathbf{M}_{i,j}=$ $~\mathbf{Y}_{1,j}+\sum_{k>1:\mathbf{Y}_{k,j}\leq\mathbf{M}_{i,j}}\mathbf{Y}_{k,j}-\mathbf{Y}_{k-1,j}$
$=$ $~\mathbf{N}_{i,jT}\mathbf{Y}_{1,j}+\sum_{k=2}^{T}\mathbf{N}_{i,j\cdot T+k}(\mathbf{Y}_{k,j}-\mathbf{Y}_{k-1,j})$
In particular, we can compute $\mathbf{M}v=\mathbf{N}v^{\prime}$ where
$v^{\prime}_{j\cdot T}=v_{j}\cdot\mathbf{Y}_{1,j}$
and
$v^{\prime}_{j\cdot T+k}=v_{j}\cdot(\mathbf{Y}_{k,j}-\mathbf{Y}_{k-1,j})\text{ for $k\geq 2$}.$
Constructing $v^{\prime}$ takes $O(Tn)$ time per query. ∎
Matrices with small constant Pollard pseudodimensions are also prevalent in real-world data, and capture a similar notion of structure as matrices with low constant VC-dimensions. For instance, the following are examples of matrices with small Pollard pseudodimension.
Definition B.12 (Multiball vectors).
Let $v\in V$ be a vertex of the graph $G$. Let $r$ be a real number and $\Delta$ be a set of distances $\infty=\delta_{0}<\delta_{1}<\dots<\delta_{\ell-1}<\delta_{\ell}=\infty$. Define the multiball vectors as follows:
[…]
@@
Theorem B.13 (Karczmarz and Zheng [2024]).
@@
For a $K_{h}$-minor-free graph $G$, for any set $\Delta$ of $\ell$ distances, the set of multiball vectors $\vec{\mathrm{MB}}_{G,\Delta}$ has pseudodimension at most h1h-1.
Appendix C Applications
@@
In this section, we consider some applications of our OMv result in Theorem˜1.4. These applications have quadratic $\Omega(n^{2})$-time lower bounds, conditional on the OMv conjecture. We show that on structured graph this lower bound can be beaten.
C.1 High-accuracy Dynamic Laplacian Solver
See 1.6
@@ C.1 High-accuracy Dynamic Laplacian Solver
Definition C.1 ($(1+\epsilon)$-spectral sparsifiers).
@@
Then, a result from Abraham et al. [2016] provides a construction for a dynamic spectral sparsifier under edge deletions and insertions with polylogarithmic amortized update time:
Theorem C.2 (Theorem 4.1 of Abraham et al. [2016]).
@@
Lemma C.3 (Kyng and Sachdeva [2016]).
@@
Via Theorem˜1.4 we have $U(n)=\widetilde{O}(n)$ and $Q(n)=\widetilde{O}(n^{2-1/d})$, thus obtaining Theorem˜1.6.
Proof.
We run two separate data structures. The first is a dynamic spectral sparsifier Theorem˜C.2 which maintains a spectral sparsifier $H\approx\mathbf{L}_{G}$ with error $\epsilon=1/2$. We also run the dynamic OMv data structure for matrix $\mathbf{A}$ (adjacency matrix).
@@
When given a vector $b\in\mathbb{R}^{V}$ and error parameter $\epsilon>0$, we use $\mathbf{H}^{-1}$ as preconditioner for the linear system $\mathbf{L}_{G}x=b$. We can multiply with $\mathbf{H}^{-1}$ by running a static Laplacian system solver Lemma˜C.3 which takes $\widetilde{O}(n)$ time per product as $\mathbf{H}$ has $\widetilde{O}(n)$ edges. In particular, we let
@@
C.2 Effective Resistance
The following Theorem˜1.7 is a corollary of the dynamic Laplacian solver from Theorem˜1.6.
See 1.7
Proof.
@@ C.2 Effective Resistance
This can be computed via the dynamic Laplacian system solver in Theorem˜1.6. ∎
C.3 Triangle Detection
See 1.8
The result follows form the following lemma for $U(n)=\widetilde{O}(n)$ and $Q(n)=\widetilde{O}(n^{2-1/d})$ from Theorem˜1.4.
Lemma C.5.
@@ C.3 Triangle Detection
C.4 Single-Source Shortest Paths and k-Center
Lemma C.6 (Brand et al. [2022]).
@@ C.4 Single-Source Shortest Paths and k-Center
A query operation which for any source node, given during the query, returns the $d$-bounded single source distances in $O(Q(n,d))$. That is, return for each $v\in V$ the distance if it is at most $d$, or $\infty$ if the distance is >d>d.
@@
See 1.9
Proof.
We first describe how to obtain a dynamic distance oracle that yields distances up to $1/\epsilon$. This is done by running the dynamic OMv data structure of Theorem˜1.4 on the adjacency matrix of $G$. Then single source distances up to $1/\epsilon$ from source vertex $u$ can be computed by repeatedly multiplying $w\leftarrow\mathbf{A}w$ for initial $w=e_{u}$, i.e., we compute $\mathbf{A}^{i}e_{u}$ for $i=1,2,...,1/\epsilon$. The smallest $i$ with $(\mathbf{A}^{i}e_{u})_{v}\neq 0$ is the smallest number of steps to reach from $u$ to $v$, i.e., the distance between the vertices.
This data structure internally constructs the a $\Delta$-labeled spanning tree of the boolean matrix $\mathbf{A}$, where the weight of the tree gives us the time complexity per query. Hence, while we do not know the corrupted VC-dimension $d$, we do know how much time each query is going to take.
So when running Lemma˜C.6 on the above distance data structure, we can pick $k=n/\sqrt{T}$ where $T\geq n$ is the weight of the spanning tree. Thus, we get the update time
@@
which proves the claim. ∎
See 1.10
This result is a corollary of Theorem˜1.9, as $k$-center can be reduced to computing $k$-single source distances.
Lemma C.7 (Theorem 7.3 by Cruciani et al. [2024]).
@@
Appendix D Characterization of matrices with constant VC-dimension
In this section, we provide a characterization of matrices with constant VC-dimension. Particularly, we prove Theorem˜1.1 and state a number of examples of other constant VC-dimension matrices.